# An Optimism-based Approach to Online Evaluation of Generative Models

### v1

@@ B.3 Optimistic marginal class distribution
B.4 Data-dependent optimistic marginal class distribution
Lemma B.2 (Optimistic marginal class distribution).
satisfies that ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})\geq H(Y_{g})$.
Proof.
It suffices that show that $\bm{\epsilon}_{g}^{n}[j]\geq|\widehat{p}_{Y_{g}}^{n}[j]-p_{Y_{g}}[j]|$ with probability at least $1-\frac{\delta}{2}$. We evoke Theorem E.4 and conclude the proof. ∎
Appendix C Additional Experimental Results
C.1 Results of FID-Based Online Evaluation
1. Embeddings Extracted by DINOv2 [33].
We run FID-UCB, naive UCB, and the greedy algorithm where the feature of an image is extracted by DINOv2-ViT-L/14. The pretrained weights are downloaded from the official repository (licensed under Apache License 2.0). [^3] https://github.com/facebookresearch/dinov2/tree/main The results are summarized in Figure 5.
CIFAR10 (G=5)
ImageNet (G=5)
FFHQ (G=5)
[…]
Figure 5: Performance of FID-UCB, naive UCB, and the greedy algorithm on various datasets, where the last row corresponds to the evaluation of the variance-controlled generators. The embeddings of the images are extracted by the pretrained DINOv2-ViT-L/14 model.
2. Embeddings Extracted by CLIP [7].
We run FID-UCB, naive UCB, and the greedy algorithm where the feature of an image is extracted by CLIP. The pretrained weights are downloaded from the official repository. [^4] https://github.com/mlfoundations/open_clip. Copyright (c) 2012-2021 Gabriel Ilharco, Mitchell Wortsman, Nicholas Carlini, Rohan Taori, Achal Dave, Vaishaal Shankar, John Miller, Hongseok Namkoong, Hannaneh Hajishirzi, Ali Farhadi, Ludwig Schmidt. The results are summarized in Figure 6. We note that in the FFHQ dataset, FID-UCB attains a low optimal pick ratio as the optimal generator and the second-best one attains almost the same CLIP-based FID.
CIFAR10 (G=5)
ImageNet (G=5)
FFHQ (G=5)
[…]
Figure 6: Performance of FID-UCB, naive UCB, and the greedy algorithm on various datasets, where the last row corresponds to the evaluation of the variance-controlled generators. The embeddings of images are extracted by the pretrained CLIP model.
3. Results on AFHQ Dog dataset.
The results are summarized in Figure 7.
AFHQ Dog (G=20)
Figure 7: Performance of FID-UCB, naive UCB, and the greedy algorithm for evaluating the variance-controlled generators on AFHQ Dog dataset. The embeddings of images are extracted by InceptionNet.V3.
C.2 Results of IS-Based Online Evaluation
1. Diversity and quality of the generated samples.
We examine the diversity and quality of the generated samples queried by IS-UCB, naive UCB, and the greedy algorithm (i.e., experiment in Figure 4). The results are summarized in Figure 8.
Step 500
[…]
Figure 8: Generated images queried by IS-UCB, naive UCB, and the greedy algorithm.
2. Results on AFHQ Dog dataset.
The results are summarized in Figure 9.
AFHQ Dog (G=20)
Figure 9: Performance of IS-UCB, naive UCB, and the greedy algorithm for evaluating the variance-controlled generators on AFHQ dataset.
Appendix D Additional Experimental details
1. Additional experiment setups.
All the experiments are run on a single Nvidia GeForce RTX 3090 GPU. For FID-based evaluation, each trial takes at most 2 hours to finish, depending on the dataset and the embedding pretrained model. For IS-based evaluation, each trial takes less than 30 minutes to finish.
2. List of pretrained generative models/generated image data.
For the CIFAR10 dataset, we compare pretrained models including SNGAN, ACGAN, BigGAN, ContraGAN, and BigGAN-CR. For the ImageNet dataset, we compare pretrained models including StyleGAN3, SNGAN, BigGAN, ContraGAN, and ReACGAN. The pretrained weights are downloaded from the StudioGAN repository [22]. For the FFHQ dataset, we compare StyleSwin, StyleNAT, StyleGAN2-ADA, StyleGAN-XL, and LDM, where we utilize the generated image datasets downloaded from the dgm-eval repository [39].
3. Implementation details of naive UCB for FID-based evaluation.
In FID-based evaluation, naive UCB upper bounds the error in estimating the mean vector by
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}=\sqrt{\sum_{j=1}^{d}(\widehat{\mu}_{g}^{n}[j]-\mu_{g}[j])^{2}}\leq\sigma_{g}\sqrt{\frac{d}{n}\log\left(\frac{2d}{\delta}\right)}$
which holds by the assumption of FID score that $f(X_{g})\sim{\mathcal{N}}(\mu_{g},\Sigma_{g})$, with probability at least $1-\delta$. In addition, the error in estimating the covariance matrix is bounded by
$\|\widehat{\Sigma}_{g}^{n}-\Sigma_{g}\|_{2}\leq\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}+\|\Sigma_{g}\|_{2}\left(2\sqrt{\frac{d}{n}}+2\sqrt{\frac{1}{2n}\log\left(\frac{2}{\delta}\right)}+\left(\sqrt{\frac{d}{n}}+\sqrt{\frac{1}{2n}\log\left(\frac{2}{\delta}\right)}\right)^{2}\right)$
which follows from the exact same derivation in Inequality (28) and holds with probability at least $1-\delta$ (see Lemma E.6). Compared to the generator-dependent error bounds utilized by FID-UCB (i.e., Inequalities (27) and (29)), these bounds are dimension-dependent and do not fully exploit the properties of the generator. Finally, by the exact same analysis in the proof of Theorem 5.1, the following estimated FID
$\widehat{\textup{FID}}_{g}^{n}=$ $\widetilde{\textup{FID}}_{g}^{n}-C_{g}^{n}\cdot\sigma_{g}\sqrt{\frac{d}{n}\log\left(\frac{6d}{\delta}\right)}-\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)}$ (32)
$-\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\left(\|\Sigma_{g}\|_{2}\left(2\sqrt{\frac{d}{n}}+2\sqrt{\frac{1}{2n}\log\left(\frac{6}{\delta}\right)}+\left(\sqrt{\frac{d}{n}}+\sqrt{\frac{1}{2n}\log\left(\frac{6}{\delta}\right)}\right)^{2}\right)+\frac{d\sigma_{g}^{2}}{n}\log\left(\frac{6d}{\delta}\right)\right)}$
lower bounds $\textup{FID}_{g}$ with probability at least $1-\delta$.
4. Implementation details of naive UCB for IS-based evaluation.
In IS-based evaluation, naive UCB uses Hoeffding-type bonus terms. Specifically, the error vector in defining the optimistic marginal class distribution is given by $e_{g}^{n}$, whose $j$-th element is defined as
$e_{g}^{n}[j]:=\sqrt{\frac{1}{2n}\log\left(\frac{4d}{\delta}\right)},$
which ensures that ${\mathcal{E}}(\textup{Clip}_{e^{-1}}(\widetilde{p}_{Y_{g}}^{n},e_{g}^{n}))\geq H(Y_{g})$ with probability at least $1-\frac{\delta}{2}$. In addition, the concentration error of $\widehat{H}^{n}(Y_{g}|X_{g})$ is bounded by
$H_{g}(Y_{g}|X_{g})\geq\widehat{H}^{n}(Y_{g}|X_{g})-C^{\prime}\sqrt{\frac{1}{2n}\log\left(\frac{4}{\delta}\right)}$
which holds with probability at least $1-\frac{\delta}{2}$. Therefore, the estimated IS is given by
$\widehat{\textup{IS}}_{g}^{n}:=\exp\left\{{\mathcal{E}}\left(\textup{Clip}_{e^{-1}}\left(\widetilde{p}_{Y_{g}}^{n},e_{g}^{n}\right)\right)-\widehat{H}^{n}(Y_{g}|X_{g})+C^{\prime}\sqrt{\frac{1}{2n}\log\left(\frac{4}{\delta}\right)}\right\}$ (33)
By the exact same analysis in the proof of Theorem 6.1, it can be shown that $\widehat{\textup{IS}}_{g}^{n}\geq\textup{IS}_{g}$ with probability at least $1-\delta$.
5. Additional implementation details of FID-UCB
To compute the bonus term (7), we need to estimate the following parameters: 1) $\|\mu_{g}-\mu_{\textup{r}}\|_{2}$: we set it to be a constant, e.g., 0.5; 2) ${\mathcal{I}}_{g}$: we estimate this term using the truncated covariance matrix with parameter $\tau$, which is denoted by $\textup{truncate}_{\tau}(\widehat{\Sigma}_{g}^{n})$. The $(i,j)$-th element of the truncated covariance matrix is $\widehat{\Sigma}_{g}^{n}[i,j]$ if $\widehat{\Sigma}_{g}^{n}[i,j]\geq\tau$ and is zero otherwise. In the experiments, we observe that a truncation parameter that is $5\sim 10\%$ of the operator norm $\|\Sigma_{\textup{r}}\|_{2}$, i.e., $0.05\cdot\|\Sigma_{\textup{r}}\|_{2}\leq\tau\leq 0.1\cdot\|\Sigma_{\textup{r}}\|_{2}$, attains satisfying performance; 3) $\|\Sigma_{g}\|_{2}$: we estimate this term by $\max_{j\in[d]}(\widehat{\sigma}_{j}^{n})^{2}+O(\sqrt{\log(1/\delta)/n})$, where $(\widehat{\sigma}_{j}^{n})^{2}$ is the sample variance for the $j$-th entry of the embedding $f(X_{g})$. This avoids the computation for the eigendecomposition of the sample covariance matrix $\widehat{\Sigma}_{g}^{n}$. The intuition is that when $\Sigma_{g}$ is diagonal, the operator norm $\|\Sigma_{g}\|_{2}$ coincides with the largest per-entry variance; 4) $\textup{Tr}[\Sigma_{g}]$: while a direct approach is to estimate it by $\textup{Tr}[\widehat{\Sigma}_{g}^{n}]$, we found that a better performance can be attained by using $O(\textup{Tr}[\textup{truncate}_{\tau}(\widehat{\Sigma}_{g}^{n})\cdot\Sigma_{\textup{r}}])$, where we utilize the fact that $\|\Sigma_{\textup{r}}^{-1}\|_{2}\cdot\textup{Tr}[\Sigma_{g}\Sigma_{\textup{r}}]\geq\textup{Tr}[\Sigma_{g}]\geq\frac{\textup{Tr}[\Sigma_{g}\Sigma_{\textup{r}}]}{\|\Sigma_{\textup{r}}\|_{2}}$.
Appendix E Auxiliary Lemmas
Lemma E.1 (Concentration of Gaussian random variable).
Let $X_{1},\cdots,X_{n}\sim{\mathcal{N}}(\mu,\sigma^{2})$. Then, it holds that
${\mathbb{P}}\left(\left|\frac{1}{n}\sum_{i=1}^{n}X_{i}-\mu\right|>\epsilon\right)<\exp\left(-\frac{n\epsilon^{2}}{2\sigma^{2}}\right)$
Equivalently, with probability at least $1-\delta$, we have that
$\left|\frac{1}{n}\sum_{i=1}^{n}X_{i}-\mu\right|\leq\sqrt{\frac{\sigma^{2}}{2n}\log\left(\frac{2}{\delta}\right)}.$
Lemma E.2 (Hoeffding’s inequality).
Let $Z,Z_{1},\cdots,Z_{n}$ be i.i.d. random variables with values in $[0,1]$ and let $\delta>0$. Then with probability at least $1-\delta$ in $(Z_{1},\cdots,Z_{n})$ we have that
$\mathbb{E}Z-\frac{1}{n}\sum_{i=1}^{n}Z_{i}\leq\sqrt{\frac{1}{2n}\log\left(\frac{1}{\delta}\right)}.$ (34)
A drawback of Hoeffding’s inequality is that the confidence interval is distribution-agnostic, and always scales with order $O(\sqrt{1/n})$. For distribution with small variance, the following [18, Bennett’s inequality] provides a tighter concentration error.
Lemma E.3 (Bennett’s inequality).
Under the conditions of Theorem E.2, with probability at least $1-\delta$ in $(Z_{1},\cdots,Z_{n})$ we have that
$\mathbb{E}Z-\frac{1}{n}\sum_{i=1}^{n}Z_{i}\leq\sqrt{\frac{2{\mathbb{V}}Z}{n}\log\left(\frac{1}{\delta}\right)}+\frac{1}{3n}\log\left(\frac{1}{\delta}\right),$ (35)
where ${\mathbb{V}}Z$ is the variance ${\mathbb{V}}=\mathbb{E}(Z-\mathbb{E}Z)^{2}$.
When the exact variance is unknown, [31] provide a purely data-dependent bound with similar properties as Bennett’s inequality.
Lemma E.4 (Restatement of [31, Theorem 4]).
Under the conditions of Theorem E.2, we have that with probability at least $1-\delta$ in the i.i.d vector $\bm{Z}=(Z_{1},\cdots,Z_{n})$ that
@@ Appendix E Auxiliary Lemmas
Lemma E.5 (Concentration of Chi-square variables).
Let $Z_{1},\cdots,Z_{n}\sim{\mathcal{N}}(0,1)$ be i.i.d standard normal. Then, it holds that
${\mathbb{P}}\left(\left|\frac{1}{n}\sum_{i=1}^{n}Z_{i}^{2}-1\right|>\epsilon\right)\leq 2e^{-n\epsilon^{2}/8}.$
Equivalently, with probability at least $1-\delta$, we have that
$\left|\frac{1}{n}\sum_{i=1}^{n}Z_{i}^{2}-1\right|\leq\sqrt{\frac{8}{n}\log\left(\frac{2}{\delta}\right)}$
Lemma E.6 (Restatement of [43, Example 6.3]).
Let $\bm{X}\in{\mathbb{R}}^{n\times d}$ be a random matrix from the $\Sigma$-Gaussian ensemble, i.e., each row vector $x_{i}$ is drawn i.i.d. from a multivariate ${\mathcal{N}}(0,\Sigma)$ distribution. Then, with probability at least $1-\delta$, we have that
$\|\widehat{\Sigma}-\Sigma\|_{2}\leq\|\Sigma\|_{2}\left(2\sqrt{\frac{d}{n}}+2\epsilon+\left(\sqrt{\frac{d}{n}}+\epsilon\right)^{2}\right),\textup{ where }\epsilon=\sqrt{\frac{1}{2n}\log\left(\frac{2}{\delta}\right)},$ (36)
where $\widehat{\Sigma}=\frac{1}{n}\sum_{i=1}^{n}x_{i}x_{i}^{\top}$ is the sample covariance matrix.
Lemma E.7 (Dimension-free concentration of sample covariance matrix [30, 25, 46]).
Let $X_{1},\cdots,X_{n}$ are i.i.d zero-mean and sub-Gaussian random vectors with parameter at most $\sigma^{2}$, whose covariance matrix is given by $\Sigma$. Let
$\bm{r}(\Sigma):=\frac{\textup{Tr}(\Sigma)}{\|\Sigma\|_{2}}$ (37)
$\left\|\frac{1}{n}\sum_{i=1}^{n}X_{i}X_{i}^{\top}-\Sigma\right\|_{2}\leq 20\sigma^{2}\|\Sigma\|_{2}\sqrt{\frac{4\bm{r}(\Sigma)+\log(1/\delta)}{n}}$ (38)
whenever $n\geq 4\bm{r}(\Sigma)+\log(1/\delta)$.
Lemma E.8.
Proof.
Lemma E.9.
Proof.
$\leq$ $\textup{Tr}[\sqrt{A}]+\textup{Tr}[\sqrt{B}],$
where the last inequality holds by $A^{-\frac{1}{2}}\succeq(A+B)^{-\frac{1}{2}}$ and $B^{-\frac{1}{2}}\succeq(A+B)^{-\frac{1}{2}}$ (see Lemma E.12). For general PSD matrices $A$ and $B$, note that
Lemma E.10.
Let $z_{1},\cdots,z_{n}\sim{\mathcal{N}}(0,1)$ are i.i.d standard Gaussian. Then, with probability at least $1-\delta$, it holds that
$\left(\frac{1}{n}\sum_{i=1}^{n}z_{i}\right)^{2}\leq\sqrt{\frac{32}{n^{2}}\log\left(\frac{2}{\delta}\right)}.$
Proof.
Note that $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}z_{i}\sim{\mathcal{N}}(0,1)$. By Lemma E.5, we have that
${\mathbb{P}}\left(\left|\left(\frac{1}{n}\sum_{i=1}^{n}z_{i}\right)^{2}-\frac{1}{n}\right|>\epsilon\right)=$ ${\mathbb{P}}\left(|U^{2}-1|>n\epsilon\right)\leq 2\exp\left(-\frac{n^{2}\epsilon^{2}}{8}\right)$
where $U^{2}\sim\chi^{2}_{1}$ is a chi-square random variable. Hence, with probability at least $1-\delta$, it holds that
$\left|\left(\frac{1}{n}\sum_{i=1}^{n}z_{i}\right)^{2}-\frac{1}{n}\right|\leq\sqrt{\frac{8}{n^{2}}\log\left(\frac{2}{\delta}\right)},$
which concludes the proof. ∎
Lemma E.11.
Let $z_{1},\cdots,z_{n},u_{1},\cdots,u_{n}\sim{\mathcal{N}}(0,1)$ be i.i.d standard Gaussian. Then, with probability at least $1-\delta$, it holds that
$\left|\left(\frac{\sum_{i=1}^{n}z_{i}}{n}\right)\left(\frac{\sum_{i=1}^{n}u_{i}}{n}\right)\right|<\frac{1}{n}\sqrt{8\log\left(\frac{4}{\delta}\right)}.$
Proof.
Note that $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}z_{i}\sim{\mathcal{N}}(0,1)$.
$\left(\frac{\sum_{i=1}^{n}z_{i}}{n}\right)\left(\frac{\sum_{i=1}^{n}u_{i}}{n}\right)=\frac{1}{n}\left(\frac{\sum_{i=1}^{n}z_{i}}{\sqrt{n}}\right)\left(\frac{\sum_{i=1}^{n}u_{i}}{\sqrt{n}}\right)=\frac{1}{n}ZU,$
where $Z,U\sim{\mathcal{N}}(0,1)$ are i.i.d standard Gaussian. Observe that
$ZU=\frac{1}{4}(Z+U)^{2}-\frac{1}{4}(Z-U)^{2}\sim\frac{1}{2}Q-\frac{1}{2}R,$
where $Q:=\frac{1}{2}(Z+U)^{2},R:=\frac{1}{2}(Z-U)^{2}\sim\chi_{1}^{2}$ are i.i.d Chi-square distributed with 1 degree of freedom (since $\frac{1}{\sqrt{2}}(Z+U)$ and $\frac{1}{\sqrt{2}}(Z-U)$ are i.i.d standard Gaussian). Hence, by Lemma E.5 and a union bound, with probability at least $1-\delta$,
$\left|\left(\frac{\sum_{i=1}^{n}z_{i}}{n}\right)\left(\frac{\sum_{i=1}^{n}u_{i}}{n}\right)\right|=$ $\left|\frac{1}{2}Q-\frac{1}{2}R\right|\leq\frac{1}{2}\left|Q-1\right|+\frac{1}{2}\left|R-1\right|$
$\leq$ $\frac{1}{n}\sqrt{8\log\left(\frac{4}{\delta}\right)},$
which concludes the proof. ∎
Lemma E.12.
Proof.

### v3

@@ B.4 Auxiliary Definitions and Lemmas
Lemma 12 (Restatement of (Maurer and Pontil,, 2009, Theorem 4)).
Let $Z,Z_{1},\cdots,Z_{n}$ be i.i.d. random variables with values in $[0,1]$ and let $\delta>0$. we have that with probability at least $1-\delta$ in the i.i.d vector $\bm{Z}=(Z_{1},\cdots,Z_{n})$ that
@@
Appendix C EXPERIMENTAL DETAILS
1. List of pretrained generative models.
For the CIFAR10 dataset, we compare pretrained generative models including iDDPM-DDIM (Nichol and Dhariwal,, 2021), LOGAN (Wu et al.,, 2020), WGAN-GP (Gulrajani et al.,, 2017), NVAE (Vahdat and Kautz,, 2020), and RESFLOW (Chen et al.,, 2019). For ImageNet, we compare pretrained models including DiT-XL-2 (Peebles and Xie,, 2023), ADMG (Dhariwal and Nichol,, 2021), BigGAN (Brock et al.,, 2019), RQ-Transformer (Lee et al.,, 2022), and ADM (Dhariwal and Nichol,, 2021). For the FFHQ dataset, we compare StyleNAT (Walton et al.,, 2023), StyleGAN2-ADA (Karras et al.,, 2020), LDM (Rombach et al.,, 2022), Unleashing-Transformers (Bond-Taylor et al.,, 2021), and Efficient-vdVAE (Hazami et al.,, 2022). We utilize the generated image datasets downloaded from the dgm-eval repository (Stein et al.,, 2023).
2. List of embeddings for FD-based evaluation and selection.
We consider three standard encoders for image data: InceptionV3.Net (Szegedy et al.,, 2016), DINOv2 (Oquab et al.,, 2024), and CLIP (Radford et al.,, 2021). Following (Stein et al.,, 2023), we utilize the DINOv2 ViT-L/14 model and the OpenCLIP ViT-L/14 trained on DataComp-1B (Gadre et al.,, 2023). Embeddings extracted by InceptionV3.Net, DINOv2, and CLIP have $2048$, $1024$, and $1024$ dimensions, respectively. For video data, we utilize the I3D model (Carreira and Zisserman,, 2018) pretrained on the Kinetics-400 (Carreira and Zisserman,, 2018) dataset following (Unterthiner et al.,, 2019), where the logits layer with $400$ dimensions is used as the embedding. For audio data, we utilize VGGish (Hershey et al.,, 2017) following (Kilgour et al.,, 2019), where the activations from the $128$ dimensional layer prior to the final classification layer are used as the embedding.
3. Implementation details of FD-UCB.
As we consider InceptionV3, DINOv2, and (unnormalized) CLIP embeddings, which are generally unbounded, we utilize the collected data to estimate the model-dependent parameters in the bonus (6). This approach preserves the online format of the FD-UCB algorithm. Our numerical results indicate that with only 50 samples, the norm terms in the bound would be close to their underlying values (Tables 12, and 3). To have a better estimation of the covariance matrix, we also adopt the thresholding method in (Cai and Liu,, 2011). We treat the batch size, parameter $\kappa$ in the bonus (6), and parameter $M$ for thresholding as hyperparameters. We conduct ablation study on these hyperparameters summarize the results in Figure 14.
Parameters $\textup{Tr}(\Sigma)$ $\textup{Tr}(\widehat{\Sigma})$ $\sqrt{\textup{Tr}[\Sigma^{2}]}$ $\sqrt{\textup{Tr}[\widehat{\Sigma}^{2}]}$ $\|\Sigma\|_{2}$ $\|\widehat{\Sigma}\|_{2}$ $\bm{r}(\Sigma)$ $\bm{r}(\widehat{\Sigma})$ Range of $\|f(X_{g})\|_{2}$
DiT-XL-2 $166.4$ $160.0$ $12.5$ $20.9$ $9.6$ $11.2$ $17.3$ $14.3$ $[9.7,33.6]$
ADMG $157.7$ $150.8$ $12.1$ $19.4$ $9.2$ $10.5$ $17.2$ $14.4$ $[10.8,34.5]$
BigGAN $154.5$ $161.8$ $13.5$ $22.3$ $11.0$ $13.2$ $14.0$ $12.2$ $[9.7,31.8]$
RQ-Transformer $166.9$ $164.5$ $13.8$ $22.2$ $10.7$ $12.7$ $15.5$ $13.0$ $[11.0,29.7]$
ADM $176.5$ $182.5$ $15.7$ $25.6$ $12.9$ $15.2$ $13.7$ $12.0$ $[0.5,34.8]$
ImageNet $181.2$ $141.2$ $12.4$ $20.2$ $9.2$ $8.7$ $19.7$ $16.2$ $[10.9,33.0]$
Table 1: Data-dependent parameters on the ImageNet dataset and standard generative models: We present both the estimated and reference values, which are computed from $50$ and $5,000$ images, respectively. The image data embeddings are extracted by InceptionV3.Net.
Parameters $\textup{Tr}(\Sigma)$ $\textup{Tr}(\widehat{\Sigma})$ $\sqrt{\textup{Tr}[\Sigma^{2}]}$ $\sqrt{\textup{Tr}[\widehat{\Sigma}^{2}]}$ $\|\Sigma\|_{2}$ $\|\widehat{\Sigma}\|_{2}$ $\bm{r}(\Sigma)$ $\bm{r}(\widehat{\Sigma})$ Range of $\|f(X_{g})\|_{2}$
DiT-XL-2 $591.2$ $601.9$ $59.1$ $101.4$ $38.2$ $48.0$ $15.5$ $12.5$ $[34.6,36.9]$
ADMG $579.9$ $566.1$ $62.1$ $101.3$ $41.2$ $50.4$ $14.1$ $11.2$ $[34.9,37.0]$
BigGAN $514.9$ $486.8$ $62.7$ $90.8$ $43.8$ $54.0$ $11.8$ $9.0$ $[34.9,36.8]$
RQ-Transformer $545.5$ $551.7$ $58.8$ $99.1$ $39.3$ $51.0$ $13.9$ $10.8$ $[35.0,36.8]$
ADM $560.2$ $561.8$ $59.0$ $100.7$ $38.3$ $52.5$ $14.6$ $10.7$ $[34.9,36.8]$
ImageNet $636.8$ $573.1$ $53.5$ $118.5$ $32.1$ $74.1$ $19.9$ $7.7$ $[35.1,36.7]$
Table 2: Data-dependent parameters on the ImageNet dataset and standard generative models: We present both the estimated and reference values, which are computed from $50$ and $5,000$ images, respectively. The image data embeddings are extracted by CLIP.
Parameters $\textup{Tr}(\Sigma)$ $\textup{Tr}(\widehat{\Sigma})$ $\sqrt{\textup{Tr}[\Sigma^{2}]}$ $\sqrt{\textup{Tr}[\widehat{\Sigma}^{2}]}$ $\|\Sigma\|_{2}$ $\|\widehat{\Sigma}\|_{2}$ $\bm{r}(\Sigma)$ $\bm{r}(\widehat{\Sigma})$ Range of $\|f(X_{g})\|_{2}$
DiT-XL-2 $2079.5$ $2085.6$ $99.9$ $309.4$ $36.5$ $81.4$ $57.0$ $25.6$ $[38.5,48.7]$
ADMG $2068.2$ $2102.1$ $103.0$ $311.4$ $42.5$ $79.1$ $48.7$ $25.6$ $[38.6,49.0]$
BigGAN $1885.5$ $1799.6$ $120.3$ $285.4$ $68.8$ $110.4$ $27.4$ $16.3$ $[36.5,48.6]$
RQ-Transformer $1955.3$ $1968.0$ $111.4$ $302.5$ $54.4$ $91.1$ $36.0$ $21.6$ $[38.4,49.3]$
ADM $2013.3$ $1988.1$ $106.7$ $306.0$ $50.5$ $101.0$ $39.9$ $19.7$ $[39.1,49.2]$
ImageNet $2112.7$ $2095.4$ $85.3$ $347.3$ $21.0$ $104.3$ $100.7$ $20.1$ $[42.2,48.3]$
Table 3: Data-dependent parameters on the ImageNet dataset and standard generative models: We present both the estimated and reference values, which are computed from $50$ and $5,000$ images, respectively. The image data embeddings are extracted by DINOv2.
4. Implementation details of Naive-UCB.
Naive-UCB is a simplification of FD-UCB and IS-UCB which replaces the generator-dependent variables in the bonus function with data-independent and dimension-based terms. For FD-based evaluation, Naive-UCB sets $\textup{Tr}[\Sigma_{g}]=O(d),\textup{Tr}[\Sigma_{g}^{2}]=O(d)$, and $\|\Sigma_{g}\|_{2}=O(1)$ in FD-UCB. For IS-based evaluation, the Naive-UCB method sets $\widehat{V}^{n}(Y_{g}|X_{g})=(\log d)^{2}$ and $\widehat{V}^{n}(p_{Y|X_{g}}[j])=1$ for any $j\in[d]$.
Avg. Regret
Figure 7: Online FD-based evaluation and selection among standard generative models: The image data embeddings are extracted by InceptionV3.Net. Results are averaged over 20 trials.
[…]
Avg. Regret
Figure 8: Online FD-based evaluation and selection among standard generative models: The image data embeddings are extracted by DINOv2-ViT-L/14. Results are averaged over 20 trials.
[…]
FFHQ (VC)
[…]
Figure 9: Online FD-based evaluation and selection among variance-controlled (VC) models: The image data embeddings are extracted by InceptionV3.Net. Results are averaged over 20 trials.
FFHQ (VC)
[…]
Figure 10: Online FD-based evaluation and selection among variance-controlled (VC) models: The image data embeddings are extracted by DINOv2. Results are averaged over 20 trials.
FFHQ (VC)
[…]
Figure 11: Online FD-based evaluation and selection among variance-controlled (VC) models: The image data embeddings are extracted by CLIP. Results are averaged over 20 trials.
MSR-VTT (Video)
[…]
Figure 12: Online evaluation and selection on video and audio data: The video data embeddings are extracted by I3D, and the audio data embeddings are extracted by VGGish. Results are averaged over 20 trials.
Figure 13: Online IS-based evaluation and selection among variance-controlled models on the AFHQ Dog dataset: IS-UCB can identify models that generate images with more diversity. Results are averaged over 20 trials.
Batch Size
Figure 14: Ablation study on the hyperparameters of FD-UCB on the CIFAR10 dataset: The image data embeddings are extracted by CLIP. Results are averaged over 20 trials.
Parameter $\kappa$
[…]
Magnatagatune (Audio)
[…]
Figure 15: Adaptation to new models: A new model is introduced after each $200$ steps. Results are averaged over 20 trials.